OpenStudy (anonymous):

Solve 3x-2y=10 2x+3y=-2

OpenStudy (anonymous):

somebody help please.

OpenStudy (anonymous):

plase help

OpenStudy (anonymous):

First try to cancel out any one variable by either addition or subtraction. Let 3x-2y=10 be eqn(1) and 2x+3y=-2 be eqn(2) To cancel out x, let's multiply eqn (1) with 2 and eqn (2) with 3 2(3x-2y)=2(10) and 3(2x+3y)=3(-2) 6x-4y=20 and 6x+9y=-6 subtract eqn(2) from eqn(1) 6x-4y=20 6x+9y=-6 (-) --------------- 0 -13y=26 --------------- -13y=26 So, y=-2 Substituting y=-2 in eqn (1), 3x-2(-2)=10 3x+4=10 x=2 So, x=2 and y=-2